OpenStudy (anonymous):

Line M is represented by the following equation: x − y = 8 Which equation completes the system that is satisfied by the solution (18, 10)? 2x − y = 26 x + y = 18 2x − 2y = 36 x − y = −28

OpenStudy (anonymous):

2x − y = 26

OpenStudy (anonymous):

why would i be that one

OpenStudy (anonymous):

plug (18,10) into each equation until you get an answer of the form x = x for example, 2x-2y=36 2(18)-2(10)=36 36-20=36 16=36 so dats not the answer

OpenStudy (anonymous):

so its not the 1st one

OpenStudy (anonymous):

Refer to the attachment from Mathematica v9.

OpenStudy (anonymous):

it won't open for me is the answer the first one

OpenStudy (anonymous):

@statefarmishere The answer is the first one

OpenStudy (anonymous):

@statefarmishere its da first 1 2x − y = 26

OpenStudy (anonymous):

thanks can you help me with another one

OpenStudy (anonymous):

Line A is represented by the following equation: x + y = 2 What is most likely the equation for line B so the set of equations has no solution? x + 2y = 2 2x + 2y = 4 2x + y = 2 x + y = 4

OpenStudy (anonymous):

if \(x+y=2\) it is not possible for \(x+y\) to also equal 4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is \[x+y=4\]

OpenStudy (anonymous):

ok thanks you sooooooooooooooooooooo much

OpenStudy (anonymous):

@statefarmishere LOL medal me

OpenStudy (anonymous):

2x + 2y = 4 if the equation above is divided by 2, then the result is x + y = 2. you have essentially one line.